XYZ owns one share of stock of ABC and one share of stock of CDE

1.

 XYZ owns one share of stock of ABC and one share of stock of CDE. The total value of his holdings is 597.40. Both stocks pay annual dividends that are expected to continue forever. The expected return on ABC stock is 8.40 percent and its annual dividend is expected to remain at $4.40 forever. The expected return on CDE stock is 9.50 percent. The next dividend paid by CDE is expected to be $3.30 and all subsequent dividends are expected to grow at a constant rate. What growth rate is expected for CDE dividends? The next dividend for both firms' stocks will be paid in one year.

 2.17% (plus or minus .02 percentage points)

 1.10% (plus or minus .02 percentage points)

 3.20% (plus or minus .02 percentage points)

 2.04% (plus or minus .02 percentage points)

 None of the above is within .02 percentage points of the correct answer

2.

 XYZ has one share of stock and one bond. The total value of the two securities is $1.400. The stock pays annual dividends. The next dividend is expected to be $24.00 and paid in one year. In two years, the dividend is expected to be $14.00 and the stock is expected to be priced at 5195.00. The stock has an expected return of 18.70 percent per year. The bond has a coupon rate of 13.60 percent and a face value of 51.000; pays semi-annual coupons with the next coupon expected in 6 months: and matures in 4 years. What is the YTM of the bond?

 6.88% (plus or minus .03 percentage points

 7.07% (plus or minus .03 percentage points

 7.29% (plus or minus .03 percentage points

 6.80% (plus or minus .03 percentage points

 None of the above is within .03 percentage points of the correct answer
Answer #1

1. Market Price of CDE = Total Value of Holdings - Value of ABC

Market Price of CDE = $97.40 - (Dividend / Expected Return)

Market Price of CDE = $97.40 - (4.40 / 8.40%)

Market Price of CDE = $97.40 - 52.38 = $45.02

Growth Rate of CDE = Expected Return - (Dividend / Price)

Growth Rate of CDE = 9.50% - (3.30 / 45.02) = 2.17% Option A

2.

Value of Stock = Dividend 1 / (1 + Rate) + (Dividend 2 + Price) / (1 + Rate)^2
Value of Stock = 24 / (1 + 0.187) + (14 + 195) / (1 + 0.187)^2
Value of Stock = $168.55

Value of Bond = Holdings - Stock = $1400 - 168.55 = $1231.45

Semi annual YTM = =RATE(8,1000*13.6%/2,-1231.45,1000) = 3.44% (Excel function)

Annual YTM = 3.44% * 2 = 6.88% Option A